# Deblending the MIGHTEE-COSMOS survey with XID+: The resolved radio source counts to Jy

### v1

@@ 4 Deblending Validation with Simulations
While XID+ has been extensively and successfully applied to confused FIR and sub-millimetre data from instruments like Herschel and SCUBA-2 (e.g., Hurley2017; Pearson2017; Shirley2021; Wang2024), its application to deep radio continuum data is less established. Radio surveys present a unique set of challenges, including more complex source morphologies (e.g., core-jet-lobe structures in AGNs; Mingo2019; Hardcastle2020), which violates the point-source assumption of XID+ (see Section 3).
@@ 4.1 The Simulation Framework
Our simulations generate realistic radio sky models that capture the morphological diversity of galaxies and the basic properties of MIGHTEE. The framework, adapted from Harrison2020 and implemented in a custom Python script, uses the galsim (Rowe2015) software package to draw galaxy models into a simulated map. The properties of the source population are drawn from the Tiered Radio Extragalactic Continuum Simulation (T-RECS) catalog (Bonaldi2019; Bonaldi2023).
@@
The modeling of galaxy morphology is tailored to the source type to achieve a high degree of realism:
@@
SFGs: These are modeled using Sersic profiles (Sersic1963). The Sersic index () is set based on the galaxy’s optical classification (e.g., for disk-like, for bulge-like), while the scale radius, flux, and orientation are taken from the T-RECS catalog values.
@@
AGNs: AGN modeling depends on their radio classification. Compact AGNs (FSRQs, BL Lacs) are modeled as single, compact Gaussian profiles.
@@
The more complex Steep-Spectrum AGNs (SS-AGNs), corresponding to Fanaroff-Riley Class I (FRI) and II (FRII) sources (Fanaroff&Riley1974), are modeled with a multi-component structure. The total flux is distributed among a core, jets, and lobes, with the fractional contribution of each component depending on the Rs parameter, defined as the ratio between the total projected source size and the projected distance between the two bright hot spots, from the T-RECS catalog, which distinguishes between core- and lobe-dominated structures. The overall source size is similarly apportioned between a central region and outer structures, with the relative extent of the jets and lobes also governed by the Rs parameter to model the brighter, more extended jets of FRIs and the dominant lobes of FRIIs.
@@
To mimic the asymmetric structures observed in real radio surveys (e.g., Laing2002), we implement a bending model where each side of the AGN is assigned randomized parameters for an angular offset, a shear, and a positional shift. The distributions from which these random values are drawn are different for FRI and FRII sources, allowing for the larger, more distorted morphologies characteristic of FRI jets. The final model components are then: a compact Gaussian core; composite jets made of a central, straight exponential profile and two outer bent exponential profiles; and lobes modeled as either diffuse Gaussians (for FRIs) or sharper Moffat profiles to represent hotspots (for FRIIs), reflecting the different termination shocks in the two classes (e.g., Hardcastle2020). Each component’s final position and orientation are determined by the bending model, ensuring a morphologically diverse and realistic AGN population. Figure 4 demonstrates the morphological complexity incorporated into our simulations. The top row displays FRI sources, characterized by their core-dominated, bent-jet structures, while the bottom row shows FRII sources with distinct hotspots and lobes. While these parametric models do not capture the full physical evolution inherent in hydrodynamic simulations (e.g., Giri2024), they successfully reproduce the structural diversity, such as bent jets and asymmetric lobes, required to challenge the deblending algorithm. This level of realism is sufficient to quantify the impact of non-point-source emission on XID+ flux recovery.
@@
The primary validation tests were conducted on a simulated map configured to match the MIGHTEE low-resolution data, with a pixel scale of 1.1 arcsec and where the source models are convolved with a Gaussian PSF of 8.6 arcsec FWHM. The simulation includes sources from the T-RECS catalog down to a flux limit of 0.5 Jy, sources fainter than this limit contribute negligibly to the final map statistics given the target noise level. For these tests, we added uniform Gaussian noise with a of 1.3 Jy/beam, a value representative of the thermal noise in the deepest parts of the MIGHTEE fields (Hale2025).
@@ 4.2 Validating the Simulation
For a quantitative test, we apply a blind source-finding algorithm Python Blob Detection and Source Finder (PyBDSF; Mohan2015) to the simulated map. We use a multi-stage process similar to that for the official MIGHTEE DR1 catalogs (Hale2025) as well as for the ELAIS-N1 field from the LOFAR Two-metre Sky Survey Deep Fields Data Release 2 (Shimwell2025). This approach is necessary for deeply confused fields like those studied here as a single-pass source-finding run is insufficient because the presence of numerous faint, undetected sources elevates the measured total noise (). This artificially high noise level can prevent the detection of other faint sources that would otherwise meet the detection threshold.
The multi-stage approach consists of three main steps. For all three steps, we run PyBDSF using its default parameters, which include a 3 island threshold and a 5 detection threshold. First, an initial pass of PyBDSF is run on the original map. This initial run generates both a catalog of the sources and a corresponding residual map. Secondly, we run PyBDSF on the residual map and we obtain a more accurate map of the background total noise. In the final step, this improved noise map is used as input for a second PyBDSF run on the original. This procedure allows the source-finder to identify faint sources, leading to a more complete and accurate source catalog.
The results from this process are shown in Figure 5, which we use to validate our simulation. At bright flux densities (Jy), the recovered flux densities are in good agreement with the input values, lying tightly along the one-to-one correlation line. This holds true for all source types, including the morphologically complex Extended AGNs and Extended SFGs. However, at flux densities below Jy, and particularly for the dominant population of Compact SFGs, a systematic overestimation of flux or "flux boosting" begins to emerge and becomes increasingly severe towards lower flux densities. This is a well-documented effect in deep radio surveys (e.g., Jauncey1968; Zwart2014 and references within), and its impact on the MIGHTEE DR1 source counts was characterized and corrected for using simulations by Hale2025. The right-hand panel quantifies this trend: the running median of the flux ratio, which is close to unity for bright sources, gradually climbs, reaching a factor of three (0.5 dex) at the faintest flux levels (Jy).
@@ 4.3.1 Masking Strategy: Threshold and Type
In addition to the flux density threshold, we test the impact of the catalog type used to generate the mask. PyBDSF’s process involves first identifying ‘islands’ of contiguous emission above a given threshold, and then fitting one or more Gaussian components to model the flux within each island. From this, it produces two primary catalogs: a Gaussian component list (GAUL), which contains the parameters of every individual Gaussian component fitted to the emission islands, and a source list (SRL), which groups related Gaussian components into what the algorithm interprets as single astrophysical objects. These catalogs are known to trace different aspects of the radio source population, particularly for bright and extended sources. By testing these different configurations, we can determine which masking approach yields the most reliable deblended flux densities for the remaining faint source population.
@@ 4.3.2 Prior list Selection
The performance of XID+ is dependent on the quality of the prior catalog. We test two scenarios:
Full T-RECS Prior: A complete prior that is analogous to using the unfiltered optical/NIR catalog for the real data. It contains the true positions of all sources from the simulation that fall within the unmasked regions.
Radio-Likely Prior: An optimized prior, the simulation analog of the Radio-Likely Prior from Section 2.4, constructed by selecting all sources from the full T-RECS catalog with an intrinsic flux greater than 1 Jy. This tests how XID+ performs with a high-purity but necessarily incomplete prior list.
@@ 4.4.1 Improvement Over Blind Source Finding
4.4.2 The Impact of Prior Catalog Quality
@@ 4.4.3 The Role of the Masking Threshold
While the prior list governs the overall precision, the choice of masking threshold affects both the precision and the accuracy of the deblended flux densities, particularly for sources near the mask limit. For both prior catalogs, masking sources 50Jy (50Jy mask) consistently produces better results compared to the 250Jy mask.
@@
We also tested whether the catalog type used to generate the mask, SRL versus the GAUL, impacts the deblended flux densities. The choice between these two is not trivial, as they are known to behave differently at both the bright and faint ends of the source population. As shown by Hale2023, the SRL catalog is more physically meaningful for bright, resolved radio galaxies, while the GAUL catalog can be a better tracer of the underlying source population in the confusion-dominated faint regime.
Given these known differences, it is important to test whether the choice of masking catalog (SRL vs. GAUL) type impacts the final deblended flux densities. A direct comparison of the results shows that the deblending performance is qualitatively and quantitatively very similar regardless of the masking catalog used. The 50 Jy threshold consistently produces lower scatter and more reliable flux densities, whether it is based on the SRL or GAUL list. The reason is that the two catalogs diverge only for the very brightest, multi-component sources, which are already removed by the 50 Jy mask.
@@ 4.4.5 The optimal strategy
The Radio-Likely prior provides substantially improved precision for individual flux density estimates, while the 50 Jy mask reduces contamination from bright, complex neighbours without introducing a systematic underestimation for faint sources. Furthermore, as shown in Section 4.4.4, the choice of masking catalog (SRL vs GAUL) has a negligible impact at this threshold. We therefore adopt the combination of the Radio-Likely prior with a 50 Jy SRL-based mask for all subsequent validation.
@@ 4.5.1 Validation with Spatially Varying Noise
To test the robustness of our strategy against realistic observational noise conditions, we apply our recommended strategy, the Radio-Likely Prior and SRL 50 Jy mask, to a larger simulation that incorporates a realistic, spatially varying noise similar to the MIGHTEE-COSMOS field. We generated an analytic model by modelling the primary beam of each of the 22 MIGHTEE-COSMOS pointings as a Gaussian and summing their squared responses on a grid matching the real data (Hale2025). This creates an idealised weight map, which is then inverted to produce a theoretical RMS map ().
This process reproduces some of the features of the real observations, including the lower noise in the field center where multiple pointings overlap, and higher noise at the edges. This pre-computed RMS map is then used as input for a galsim VariableGaussianNoise generator, which adds the appropriate level of noise to each pixel of the noiseless simulated sky (Rowe2015).
@@ 4.5.2 Positional Offsets
In addition to noise variations, we tested the impact of positional offsets between radio and optical/infrared hosts, which can arise from both astrometric uncertainties and genuine physical displacements such as extended jets. To assess the robustness of our strategy to such offsets, we performed a dedicated stress test in which we shifted the positions of all simulated sources using an empirical offset distribution measured from the MIGHTEE Early Science multi-wavelength catalog (Whittam2024). These offsets were derived from 6,102 sources securely matched using the likelihood ratio method, ensuring they represent the true astrometric scatter of the survey. We applied these offsets to the input catalog before re-running our adopted configuration (Radio-Likely prior, 50 Jy mask). This process reduced the number of successfully matched sources (within 2.5 arcsec) from 67,307 to 62,878, a loss of 7 per cent, providing a realistic estimate of the catalog yield under imperfect positional assumptions.
@@ 4.6.1 Flux density and uncertainty recovery
The observed distribution is complex. The peak is shifted slightly to the left of zero, while it also shows an excess of sources with positive residuals compared to the ideal Gaussian, particularly in the 0 to +2 range. These competing features result in a global mean for the distribution that is extremely close to zero. This demonstrates that while there are competing statistical effects on individual sources, the catalog as a whole is unbiased.
@@ 4.6.2 Individual Flux density Uncertainties
To understand the nature of the individual uncertainties, we examine the asymmetry of the posterior probability distributions, as shown in Figure 11. We quantify the shape of the posterior by the ratio of the upper uncertainty (84th - 50th percentile) to the lower uncertainty (50th - 16th percentile). A ratio of 1 indicates a symmetric, Gaussian-like posterior. For the vast majority of sources, the posterior remains symmetric (ratio 1) down to flux densities comparable to, and even slightly below, the 1.3 Jy (1) noise limit. This indicates that for most sources at the noise level, the uncertainties are well-behaved. For faint sources below the noise limit, the ratio becomes . This is an expected consequence of the posterior being truncated by the physical non-negativity constraint on flux density, resulting in a skewed distribution with a long tail towards higher values. Interestingly, the plot also reveals a population of sources at intermediate flux densities that exhibit highly skewed uncertainties (ratio > 2), raising questions about the reliability of these flux measurements.
@@ 4.6.3 Assessing Detection Significance
Based on this, we provide two recommendations for using the deblended catalog. For studies that can incorporate the full posterior probability distribution for each source, the entire catalog can be used, as the posteriors correctly capture the complex uncertainties. However, for many science cases, a simple, high-purity sample is required. For these applications, we recommend a selection based on significance. A cut of Significance 3 provides a practical way to flag a reliable subset of detections. Given its usefulness for assessing detection reliability, we will provide the significance for each source in the final deblended catalog.

### v2

@@ 4 Deblending Validation with Simulations
While XID+ has been extensively and successfully applied to confused FIR and sub-millimetre data from instruments like Herschel and SCUBA-2 (e.g., Hurley et al. 2017; Pearson et al. 2017; Shirley et al. 2021; Wang et al. 2024a), its application to deep radio continuum data is less established. Radio surveys present a unique set of challenges, including more complex source morphologies (e.g., core-jet-lobe structures in AGNs; Mingo et al. 2019; Hardcastle and Croston 2020), which violates the point-source assumption of XID+ (see Section 3).
@@ 4.1 The Simulation Framework
Our simulations generate realistic radio sky models that capture the morphological diversity of galaxies and the basic properties of MIGHTEE. The framework, adapted from Harrison et al. (2020) and implemented in a custom Python script, uses the galsim (Rowe et al., 2015) software package to draw galaxy models into a simulated map. The properties of the source population are drawn from the Tiered Radio Extragalactic Continuum Simulation (T-RECS) catalogue (Bonaldi et al., 2019, 2023).
@@
The modelling of galaxy morphology is tailored to the source type to achieve a high degree of realism:
@@
SFGs: These are modeled using Sersic profiles (Sérsic, 1963). The Sersic index () is set based on the galaxy’s optical classification (e.g., for disk-like, for bulge-like), while the scale radius, flux, and orientation are taken from the T-RECS catalogue values.
@@
AGNs: AGN modelling depends on their radio classification. Compact AGNs (FSRQs, BL Lacs) are modeled as single, compact Gaussian profiles.
@@
The more complex Steep-Spectrum AGNs (SS-AGNs), corresponding to Fanaroff-Riley Class I (FRI) and II (FRII) sources (Fanaroff and Riley, 1974), are modeled with a multi-component structure. The total flux is distributed among a core, jets, and lobes, with the fractional contribution of each component depending on the Rs parameter, defined as the ratio between the total projected source size and the projected distance between the two bright hot spots, from the T-RECS catalogue, which distinguishes between core- and lobe-dominated structures. The overall source size is similarly apportioned between a central region and outer structures, with the relative extent of the jets and lobes also governed by the Rs parameter to model the brighter, more extended jets of FRIs and the dominant lobes of FRIIs.
@@
To mimic the asymmetric structures observed in real radio surveys (e.g., Laing and Bridle 2002), we implement a bending model where each side of the AGN is assigned randomized parameters for an angular offset, a shear, and a positional shift. The distributions from which these random values are drawn are different for FRI and FRII sources, allowing for the larger, more distorted morphologies characteristic of FRI jets. The final model components are then: a compact Gaussian core; composite jets made of a central, straight exponential profile and two outer bent exponential profiles; and lobes modeled as either diffuse Gaussians (for FRIs) or sharper Moffat profiles to represent hotspots (for FRIIs), reflecting the different termination shocks in the two classes (e.g., Hardcastle and Croston 2020). Each component’s final position and orientation are determined by the bending model, ensuring a morphologically diverse and realistic AGN population. Figure 4 demonstrates the morphological complexity incorporated into our simulations. The top row displays FRI sources, characterized by their core-dominated, bent-jet structures, while the bottom row shows FRII sources with distinct hotspots and lobes. While these parametric models do not capture the full physical evolution inherent in hydrodynamic simulations (e.g., Giri et al., 2024) or the full morphological diversity revealed by the sensitivity of MeerKAT, this level of realism is sufficient for our purposes. The brightest, most complex sources are masked in our analysis, meaning XID+ is primarily tasked with deblending the fainter, likely to be more compact population, which is well-represented by our simulation.
@@
The primary validation tests were conducted on a simulated map configured to match the MIGHTEE low-resolution data, with a pixel scale of 1.1 arcsec and where the source models are convolved with a Gaussian PSF of 8.6 arcsec FWHM. The simulation includes sources from the T-RECS catalogue down to a flux limit of 0.5 Jy, sources fainter than this limit contribute negligibly to the final map statistics given the target noise level. For these tests, we added uniform Gaussian noise with a of 1.3 Jy/beam, a value representative of the thermal noise in the deepest parts of the MIGHTEE fields (Hale et al., 2025).
@@ 4.2 Validating the Simulation
For a quantitative test, we apply a blind source-finding algorithm Python Blob Detection and Source Finder (PyBDSF; Mohan and Rafferty 2015) to the simulated map. We use a multi-stage process similar to that for the official MIGHTEE DR1 catalogues (Hale et al. 2025) as well as for the ELAIS-N1 field from the LOFAR Two-metre Sky Survey Deep Fields Data Release 2 (Shimwell et al., 2025). This approach is necessary for deeply confused fields like those studied here as a single-pass source-finding run is insufficient because the presence of numerous faint, undetected sources elevates the measured total noise (). This artificially high noise level can prevent the detection of other faint sources that would otherwise meet the detection threshold.
The multi-stage approach consists of three main steps. For all three steps, we run PyBDSF using its default parameters, which include a 3 island threshold and a 5 detection threshold. First, an initial pass of PyBDSF is run on the original map. This initial run generates both a catalogue of the sources and a corresponding residual map. Secondly, we run PyBDSF on the residual map and we obtain a more accurate map of the background total noise. In the final step, this improved noise map is used as input for a second PyBDSF run on the original. This procedure allows the source-finder to identify faint sources, leading to a more complete and accurate source catalogue.
The results from this process are shown in Figure 5, which we use to validate our simulation. At bright flux densities (Jy), the recovered flux densities are in good agreement with the input values, lying tightly along the one-to-one correlation line. This holds true for all source types, including the morphologically complex Extended AGNs and Extended SFGs.
However, at flux densities below Jy, and particularly for the dominant population of Compact SFGs, a systematic overestimation of flux or "flux boosting" begins to emerge and becomes increasingly severe towards lower flux densities. This is a well-documented effect in deep radio surveys (e.g., Jauncey 1968; Zwart et al. 2015 and references within), and its impact on the MIGHTEE DR1 source counts was characterized and corrected for using simulations by Hale et al. (2025). The right-hand panel quantifies this trend: the running median of the flux ratio, which is close to unity for bright sources, gradually climbs, reaching a factor of three (0.5 dex) at the faintest flux levels (Jy).
@@ 4.3.1 Masking Strategy: Threshold and Type
In addition to the flux density threshold, we test the impact of the catalogue type used to generate the mask. PyBDSF’s process involves first identifying ‘islands’ of contiguous emission above a given threshold, and then fitting one or more Gaussian components to model the flux within each island. From this, it produces two primary catalogues: a Gaussian component list (GAUL), which contains the parameters of every individual Gaussian component fitted to the emission islands, and a source list (SRL), which groups related Gaussian components into what the algorithm interprets as single astrophysical objects. These catalogues are known to trace different aspects of the radio source population, particularly for bright and extended sources. By testing these different configurations, we can determine which masking approach yields the most reliable deblended flux densities for the remaining faint source population.
@@ 4.3.2 Prior List Selection
The performance of XID+ is dependent on the quality of the prior catalogue. We test two scenarios:
Full T-RECS Prior: A complete prior that is analogous to using the unfiltered optical/NIR catalogue for the real data. It contains the true positions of all sources from the simulation that fall within the unmasked regions.
Radio-Likely Prior: An optimized prior, the simulation analog of the Radio-Likely Prior from Section 2.4, constructed by selecting all sources from the full T-RECS catalogue with an intrinsic flux greater than 1 Jy. This tests how XID+ performs with a high-purity but necessarily incomplete prior list.
@@ 4.4.1 Improvement Over Blind Source Finding
4.4.2 The Impact of Prior Catalogue Quality
@@ 4.4.3 The Role of the Masking Threshold
While the prior list governs the overall precision, the choice of masking threshold affects both the precision and the accuracy of the deblended flux densities, particularly for sources near the mask limit. For both prior catalogues, masking sources 50Jy (50Jy mask) consistently produces better results compared to the 250Jy mask.
@@
We also tested whether the catalogue type used to generate the mask, SRL versus the GAUL, impacts the deblended flux densities. The choice between these two is not trivial, as they are known to behave differently at both the bright and faint ends of the source population. As shown by Hale et al. (2022), the SRL catalogue is more physically meaningful for bright, resolved radio galaxies, while the GAUL catalogue can be a better tracer of the underlying source population in the confusion-dominated faint regime.
Given these known differences, it is important to test whether the choice of masking catalogue (SRL vs. GAUL) type impacts the final deblended flux densities. A direct comparison of the results shows that the deblending performance is qualitatively and quantitatively very similar regardless of the masking catalogue used. The 50 Jy threshold consistently produces lower scatter and more reliable flux densities, whether it is based on the SRL or GAUL list. The reason is that the two catalogues diverge only for the very brightest, multi-component sources, which are already removed by the 50 Jy mask.
@@ 4.4.5 The Optimal Strategy
The Radio-Likely prior provides substantially improved precision for individual flux density estimates, while the 50 Jy mask reduces contamination from bright, complex neighbours without introducing a systematic underestimation for faint sources. Furthermore, as shown in Section 4.4.4, the choice of masking catalogue (SRL vs GAUL) has a negligible impact at this threshold. We therefore adopt the combination of the Radio-Likely prior with a 50 Jy SRL-based mask for all subsequent validation.
@@ 4.5.1 Validation with Spatially Varying Noise
To test the robustness of our strategy against realistic observational noise conditions, we apply our recommended strategy, the Radio-Likely Prior and SRL 50 Jy mask, to a larger simulation that incorporates a realistic, spatially varying noise similar to the MIGHTEE-COSMOS field. We generated an analytic model by modelling the primary beam of each of the 22 MIGHTEE-COSMOS pointings as a Gaussian and summing their squared responses on a grid matching the real data (Hale et al., 2025). This creates an idealised weight map, which is then inverted to produce a theoretical RMS map ().
This process reproduces some of the features of the real observations, including the lower noise in the field centre where multiple pointings overlap, and higher noise at the edges. This pre-computed RMS map is then used as input for a galsim VariableGaussianNoise generator, which adds the appropriate level of noise to each pixel of the noiseless simulated sky (Rowe et al., 2015).
@@ 4.5.2 Positional Offsets
In addition to noise variations, we tested the impact of positional offsets between radio and optical/infrared hosts, which can arise from both astrometric uncertainties and genuine physical displacements such as extended jets. To assess the robustness of our strategy to such offsets, we performed a dedicated stress test in which we shifted the positions of all simulated sources using an empirical offset distribution measured from the MIGHTEE Early Science multi-wavelength catalogue (Whittam et al., 2024). These offsets were derived from 6,102 sources securely matched using the likelihood ratio method, ensuring they represent the true astrometric scatter of the survey. We applied these offsets to the input catalogue before re-running our adopted configuration (Radio-Likely prior, 50 Jy mask). This process reduced the number of successfully matched sources (within 2.5 arcsec) from 67,307 to 62,878, a loss of 7 per cent, providing a realistic estimate of the catalogue yield under imperfect positional assumptions.
@@ 4.6.1 Flux Density and Uncertainty Recovery
The observed distribution is complex. The peak is shifted slightly to the left of zero, while it also shows an excess of sources with positive residuals compared to the ideal Gaussian, particularly in the 0 to +2 range. These competing features result in a global mean for the distribution that is extremely close to zero. This demonstrates that while there are competing statistical effects on individual sources, the catalogue as a whole is unbiased.
@@ 4.6.2 Individual Flux Density Uncertainties
To understand the nature of the individual uncertainties, we examine the asymmetry of the posterior probability distributions, as shown in Figure 11. We quantify the shape of the posterior by the ratio of the upper uncertainty (84th - 50th percentile) to the lower uncertainty (50th - 16th percentile). A ratio of 1 indicates a symmetric, Gaussian-like posterior. For the vast majority of sources, the posterior remains symmetric (ratio 1) down to flux densities comparable to, and even slightly below, the 1.3 Jy/beam (1) noise limit. This indicates that for most sources at the noise level, the uncertainties are well-behaved. For faint sources below the noise limit, the ratio becomes . This is an expected consequence of the posterior being truncated by the physical non-negativity constraint on flux density, resulting in a skewed distribution with a long tail towards higher values. Interestingly, the plot also reveals a population of sources at intermediate flux densities that exhibit highly skewed uncertainties (ratio > 2). This behaviour is expected: at lower flux densities, the posterior becomes increasingly affected by the physical non-negativity constraint on flux density, leading to asymmetric uncertainties. This is a natural consequence of lower S/N rather than an indication of unreliable measurements.
@@ 4.6.3 Assessing Detection Significance
Based on this, we provide two recommendations for using the deblended catalogue. For studies that can incorporate the full posterior probability distribution for each source, the entire catalogue can be used, as the posteriors correctly capture the complex uncertainties. However, for many science cases, a simple, high-purity sample is required. For these applications, we recommend a selection based on significance. A cut of Significance 3 provides a practical way to flag a reliable subset of detections. Given its usefulness for assessing detection reliability, we will provide the significance for each source in the final deblended catalogue.